OpenStudy (anonymous):

I already did the steps..just need help if my answers are correct. Solve 3x^2 – 2x – 9 = 0 3x^2 - 2x = 9 x^2 - (2/3)x = 3 (x - 1/3)^2 = 10/9 x = (1 - sqroot10)/3 x = (1 + sqroot 10)/3

OpenStudy (anonymous):

The roots are \[ \left\{\left\{x\to \frac{1}{3} \left(1-2 \sqrt{7}\right)\right\},\left\{x\to \frac{1}{3} \left(1+2 \sqrt{7}\right)\right\}\right\} \]

OpenStudy (anonymous):

so my answers are correct or wrong?

OpenStudy (anonymous):

They are not correct.

OpenStudy (anonymous):

ok